We will use this prompt for this and the next 1 question.

A survey reported in Time magazine included the question ‘‘Do you favor a federal law requiring a 15 day waiting period to purchase a gun?” Results from a random sample of US citizens showed that 318 of the 520 men who were surveyed supported this proposed law while 379 of the 460 women sampled said ‘‘yes”.

Use this information to find a 90% confidence interval for the difference in the two proportions, pm - pw, where pm is the proportion of men who support the proposed law and pw is the proportion of women who support the proposed law.

The standard error of the statistic in this situation is 0.028

What is the lower bound of your confidence interval? Round your answer to three decimal places.

A survey reported in Time magazine included the question ‘‘Do you favor a federal law requiring a 15 day waiting period to purchase a gun?” Results from a random sample of US citizens showed that 318 of the 520 men who were surveyed supported this proposed law while 379 of the 460 women sampled said ‘‘yes”.

Use this information to find a 90% confidence interval for the difference in the two proportions, pm - pw, where pm is the proportion of men who support the proposed law and pw is the proportion of women who support the proposed law.

The standard error of the statistic in this situation is 0.028

What is the upper bound of your confidence interval? Round your answer to three decimal places.
Answer #1

Ans:

sample proportion for men=318/520=0.6115

sample proportion for women=379/460=0.8239

standard error for difference=SQRT((0.6115*(1-0.6115)/520)+(0.8239*(1-0.8239)/460))=0.028

90% confidence interval for difference in proportions

=(0.6115-0.8239)+/-1.645*0.028

=-0.212+/-0.046

=(-0.258, -0.166)